OpenStudy (anonymous):

The volume of sphere W is approximately 2.93pie in.3 If the scale factor from sphere W to sphere X is 3, what is the diameter of sphere X?

OpenStudy (campbell_st):

the volume is 27 times larger... simple proof is r = 1 and r = 3 \[V _{3} = 4/3 \pi 3^3\] \[V _{1} = 4/3 \pi 1^3\] then its \[V _{3}/V _{1} = 27\] so multiple the original volume by 27 then solve for r the double your answer

OpenStudy (campbell_st):

\[2.93 \times 27 = 4/3 \pi r^3\]